Chuck and Dana agree to meet in Chicago for the weekend. Chcuk travels 114 miles in the same time that Dana travels 96 miles. If Chuck's rate of travel is 6 mph more than Dana's, then at what rate does Chuck travel?
Answer by checkley77(12844)

D=RT
114=(R+6)T
T=114/(R+6) FOR CHUCK'S TRIP.
96=RT
T=96/R FOR DANA'S TRIP.
SEEING AS THE TIMES ARE THE SAME SET THESE 2 EQUATIONS EQUAL.
114/(R+6)=96/R CROSS MULTIPLY
96(R+6)=114R
96R+576=114R
96R-114R=-576
-18R=-576
R=-576/-18
R=32 MPH. IS DANA'S SPEED.
32+6=38 MPH. IS CHUCK'S SPEED.
PROOF:
114/38=96/32
3=3